# Multispectral State-Space Feature Fusion: Bridging Shared and Cross-Parametric Interactions for Object Detection

### v1

@@ 5 Conclusion
In this paper, we propose an MS2Fusion method for multispectral image feature fusion. It leverages dynamic state space models to efficiently integrate cross-modal features, aiming to enhance object detection accuracy while reducing computational complexity. To further enhance feature fusion, a shared-parameter state space module is introduced to extract shared features across modalities, which strengthens the representation of individual modal features and facilitates effective complementarity and communication between different modal features. Incorporating shared feature modules not only improves the efficiency and precision of feature fusion but also mitigates the challenges posed by information imbalances and differences between modalities. Comprehensive validation on multiple downstream tasks as well as multiple datasets shows that the proposed method achieves state-of-the-art performance, validating its effectiveness and superiority.

### v2

@@ 4.6.3 Visualization of Feature Fusion Comparisons
4.6.4 Visual Analysis of Shared and Complementary Features.
As illustrated in Figure 12, the first column shows the RGB and thermal input images, while the second and third columns visualize the complementary and shared feature maps learned by CP-SSM and SP-SSM, respectively. Due to the low-light conditions of this scene, the RGB modality captures relatively weak and sparse features, whereas the thermal modality provides richer structural and intensity cues. The CP-SSM focuses on these modality-specific differences, emphasizing distinctive thermal gradients and texture variations. In contrast, the SP-SSM effectively fuses information from both modalities, producing shared semantic representations that consistently highlight object contours, pedestrian regions, and road boundaries across RGB and thermal domains. This indicates that SP-SSM compensates for illumination-induced degradation in the visible spectrum by enforcing semantic consistency and cross-modal alignment. Overall, this visualization demonstrates that MS2Fusion’s dual-path design enables robust feature learning under challenging night-time conditions, balancing modality complementarity and shared semantic understanding.
Figure 12: Visualization of intermediate features under low-light conditions. The first column shows RGB and thermal inputs; the second and third columns display complementary (CP-SSM) and shared (SP-SSM) features, respectively.
@@ 5 Conclusion
In this paper, we propose MS2Fusion, an adaptive and efficient multispectral feature fusion framework that addresses the key limitations of existing methods. Unlike conventional approaches that rely on rigid fusion strategies or suffer from computational inefficiency, our method leverages dynamic state space models to efficiently capture long-range dependencies while maintaining low computational complexity, overcoming the drawbacks of both CNN-based and Transformer-based techniques. Furthermore, we introduce a shared-parameter state space module to explicitly model cross-modal feature interactions, enhancing discriminative power while mitigating information loss and modal discrepancies. Extensive experiments demonstrate that MS2Fusion achieves state-of-the-art performance, validating its effectiveness in robust multispectral object detection under challenging conditions.
In future work, we plan to adapt MS2Fusion for resource-constrained platforms such as wireless sensor networks and UAV-based systems. Leveraging recent progress in efficient data transmission [100] and lightweight real-time detection for drones [101], we aim to integrate TinyML-oriented optimizations, including model pruning, quantization, and state-space compression, to enhance real-time efficiency and energy awareness. Beyond the current scope, the proposed fusion methodology will be extended to other multimodal sensing tasks under computational and bandwidth constraints, supported by our ongoing construction of a multispectral UAV–maritime dataset for large-scale validation.